OpenStudy (anonymous):

Combining Functions to Write a New Function that Models a Real-World Situation A small publishing company is releasing a new book. The production costs will include a one-time fixed cost for editing and an additional cost for each book printed. The total production cost C (in dollars) is given by the function C=550=18.95N , where N is the number of books. The total revenue earned (in dollars) from selling the books is given by the function R=33.70N . Let P be the profit made (in dollars). Write an equation relating P to N . Simplify your answer as much as possible.

OpenStudy (jim766):

profit = Revenue - Costs this is like the last one, but in this one you subtract

OpenStudy (jim766):

what is revenue (R) equal to?

OpenStudy (anonymous):

@Jim766 how do you know what to add/subtract?

OpenStudy (jim766):

In the problem what is the revenue?

OpenStudy (anonymous):

r=33.70N

OpenStudy (jim766):

P = 33.70N - C and what is the cost, according to the problem

OpenStudy (anonymous):

c=550+18.95N

OpenStudy (jim766):

P = R - C P = 33.70N - (550 + 18.95N) I know to subtract in this case, because I know that Profit = Rev - Cost They just expect you to know that...

OpenStudy (jim766):

do you know how to subtract this, to simplify it

OpenStudy (anonymous):

@Jim766 I get that you can solve it, but I'm having difficulty knowing when to add/subtract what variables

OpenStudy (anonymous):

So what you're saying is I don't have to be a moron to get this?

OpenStudy (jim766):

you can get this...lol P = 33.70N - (550 + 18.95N) p = 33.70N -550 - 18.95N

OpenStudy (jim766):

now just combine like terms

OpenStudy (anonymous):

@Jim766 I really appreciate your help, thanks a lot

OpenStudy (jim766):

NP